Circular No. 8358

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 Q4 (NEAT) Further to IAUC 8351, W. J. Carpenter (University of Cincinnati),

M. L. Sitko, R. W. Russell, and D. L. Kim report additional 3–13-μm spectrophotometry of C/2001 Q4, obtained on June 17.2 UT at Mt. Lemmon (integration time 45 min; reference star α Boo): “A smooth comet continuum was seen to rise from 3.5 to 8.4 μm, beyond which a moderate silicate emission band was observed. An underlying blackbody continuum with a temperature of ∼ 305 ± 5 K was fit to the continuum flux at 5, 8.4, and 12 μm. This grain temperature is ∼ 16 ± 2 percent higher than that of an equilibrium blackbody at the heliocentric distance of the comet. The silicate-feature-to-continuum ratio was ∼ 1.26 ± 0.02 — lower than the values of 1.65 reported on May 14.2 (IAUC 8342) and 1.43 reported for May 31.2 (IAUC 8351). During the period of observation, the strength of the silicate feature above the continuum was unchanged. The crystalline olivine feature at 11.2 μm continues to be present with approximately the same contrast as orginally seen. With our aperture, the comet has the following narrowband (∼ 0.25 μm) magnitudes and combined random errors: [3.7 μm] = 7.75 _{−0.33}^{+0.45}, [4.7 μm] = 5.68 _{−0.17}^{+0.20}, [5 μm] = 5.61 _{−0.18}^{+0.23}, [8 μm] = 1.96 ± 0.10, [10.5 μm] = 0.43 ± 0.03, [12 μm] = 0.15 ± 0.02.”

COMETS C/2003 Y11–Y15 (SOHO) Further to IAUC 8357, measurements have been reported by K. Bat-

tams (C/2003 Y11) and K. McGleam (remainder) of additional Kreutz sungrazing comets found on SOHO website images by M. Oates (C/2003 Y11), R. Matson (2003 Y12, Y14, Y15), and X.-m. Zhou (C/2003 Y13). C/2003 Y13 and Y15 were observed in C3 images, while C/2003 Y11–Y14 were all visible on C2 images. The reductions by R. Kisala and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2003 UT  α_{2000}  δ_{2000}  MPEC
C/2003 Y11 Dec. 26.314  18^{h}24.^{m}2  −24^{o}51′  2004-L67
C/2003 Y12  27.333  18 28.4  −24 46  2004-L68
C/2003 Y13  27.351  18 36.5  −29 21  2004-L68
C/2003 Y14  27.375  18 29.1  −24 44  2004-L68
C/2003 Y15  31.870  18 53.8  −25 06  2004-L68

SUPERNOVA 2004ci IN NGC 5980

The report by Park and Foley on IAUC 8357 was an independent discovery.

2004 June 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green